SunnyLandscape04:

How do I do a 2-step equation?

Creed:

Depending on the situation, can I see the equation? Sincerely, ℂ𝕣𝕖𝕖𝕕

snowflake0531:

There are 3 ways to solve for the two variables in a two step equation 1. Graphing method 2. Substitution method 3. Elimination method Graphing method is the easiest, just graph it on your graphing calculator or desmos. The intersecting coordinate point is your x and y (x,y) The substitutionmethod: Take one equation, and solve for what x or y equals, incorporating the other variable in it tho, so you have for example x = ___ or y=___ Now that you have that, you can substitute the x or y value into the second equation for example, if x was 2y-1, and the second equation was 2y+2x=1 (completely random, may not even be an actual solution), but you would do 2y+2(2y-1)=1 As this, you can solve for y After you get y, you can substiute the actual y value, into one of the two equations, finding x The elimination method is used the least In this way,you are multiply a whole equation by a number, making it, for example we have 2x+3y=7 3x+2y=7 So we can multiply the top equation my 3, the bottom equation by 2, so that the top equation will get 6x and the bottom 6x, so we can subtract the equations together, and get rid of the x's So, all we have left is _y=___ So, by this, we can solve for y Then we can substitute the y value into one of hte two equations to solve for x

Creed:

Wow, I can tell that took a ton of time and that deserves a great amount of appreciation! Nice Job snowflake! :) Sincerely, ℂ𝕣𝕖𝕖𝕕

SunnyLandscape04:

Okay thank you.

snowflake0531:

yw~